# Studying Generalization on Memory-Based Methods in Continual Learning

### v1

@@ 3 Experimental Setting
We create a dataset where we control every characteristic $z$ that generates an image. We identify one of these characteristics as the label $y$, a group of relevant features to solve the task $z_{g}$, and the rest as irrelevant features. In order for the model to be effective, it must identify features $z_{g}$ and combine them to correctly identify $y$. We expect that even with missing combinations of $y$ $z_{g}$ in the training set, a robust model that is able to identify $y$ and $z_{g}$ independently can extrapolate its knowledge correctly to solve the task despite the absent combinations.
For the CL scenario, we create the sequence following a domain incremental setup (Van de Ven & Tolias, 2019), each experience contains every class and a different domain distribution. Each experience will have a disjoint group of features $z_{g}$ present. For every experience $t\in E$, $z_{t}=\{\hat{z}_{0},\hat{z}_{1},...,\hat{z}_{E}\}$, where $z_{t_{1}}\cap z_{t_{2}}=\emptyset,\forall{t_{1}}\neq{t_{2}}$. We create a test set that follows the same distribution $P_{t}(y_{i}|x_{i},z_{i})$ of each experience. In addition to the test set, we built a second set, which we call a generalization set, to test the OOD generalization capabilities of the model. This is achieved by holding out a number of combinations of features-label tuples $(z_{j},y_{j})\in G_{t}$ from being given to the model during training $P_{t}(z_{j},y_{j})=0,\forall t\in E$. An example of the training and testing scenario generated can be seen in Fig. 1.
@@ 6 Conclusion and future work
Memory-based methods have shown high performance in various CL scenarios. However, the generalizability of these models is rarely tested. In this work, we show that these methods can only generalize in the limited context provided by the buffer only with enough memory size, and that for OOD elements the performance is low. We believe it is important to expand these types of studies to better comprehend these techniques and then propose alternatives that can generalize out-of-distribution. In future work, we seek to propose methods that are capable to better take advantage of learned representations to increase their ability to generalize.

### v2

@@ 3 Experimental Setting
We create a dataset where we control every characteristic $z$ that generates an image. We identify one of these characteristics as the label $y$, a group of relevant features to solve the task $z_{g}$, and the rest as irrelevant features. In order for the model to be effective, it must identify features $z_{g}$ and combine them to correctly identify $y$. We expect that even with missing combinations of $y$, $z_{g}$ from the training set, a robust model that is able to identify $y$ and $z_{g}$ independently can extrapolate its knowledge correctly to solve the task despite the absent combinations.
For the CL scenario, we create the sequence following a domain incremental setup (Van de Ven & Tolias, 2019), each experience contains every class and a different domain distribution. Each experience will have a disjoint group of features $z_{g}$ present. For every experience $t,t_{1},t_{2}\in\{1,2,...,E\}$, $z_{t}=\{\hat{z}_{0},\hat{z}_{1},...,\hat{z}_{|z_{t}|}\}$, where $z_{t_{1}}\cap z_{t_{2}}=\emptyset,\forall{t_{1}}\neq{t_{2}}$. We create a test set that follows the same distribution $P_{t}(y_{i}|x_{i},z_{i})$ of each experience. In addition to the test set, we built a second set, which we call a generalization set, to test the OOD generalization capabilities of the model. This is achieved by holding out a number of combinations of features-label tuples $(z_{j},y_{j})\in G_{t}$ from being given to the model during training $P_{t}(z_{j},y_{j})=0,\forall t\in E$. An example of the training and testing scenario generated can be seen in Fig. 1.
@@ 6 Conclusion and future work
Memory-based methods have shown high performance in various CL scenarios. However, the generalizability of these models is rarely tested. In this work, we show that these methods can only generalize in the limited context provided by the buffer only with enough memory size and that for OOD elements the performance is low. We believe it is essential to expand these types of studies to better comprehend these techniques and then propose alternatives that can generalize out-of-distribution. In future work, we seek to propose new methods that are capable of better-taking advantage of learned representations to increase their ability to generalize.
7 Acknowledgement
Research partly funded by National Center for Artificial Intelligence CENIA, FB210017, BASAL, ANID, and by PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - ”FAIR - Future Artificial Intelligence Research” - Spoke 1 ”Human-centered AI”, funded by the European Commission under the NextGeneration EU programme.